# Discovery of the optical counterpart of the fast X-ray transient EP240414a

### v1

@@ Abstract
Fast X-ray transients (FXTs) are extragalactic bursts of X-rays first identified in archival X-ray data, and now routinely discovered by the Einstein Probe in real time, which is continuously surveying the night sky in the soft ($0.5-4$ keV) X-ray regime. In this Letter, we report the discovery of the second optical counterpart (AT 2024gsa) to an FXT (EP 240414a). EP 240414a is located at a projected radial separation of 27 kpc from its likely host galaxy at $z=0.4018\pm 0.0010$. The optical light curve of AT 2024gsa displays three distinct components. The initial decay from our first observation is followed by a re-brightening episode, displaying a rapid rise in luminosity to an absolute magnitude of $M_{r}\sim-21$ after two rest-frame days. While the early optical luminosity and decline rate is similar to luminous fast blue optical transients, the colour temperature of AT 2024gsa is distinctly red and we show that the peak flux is inconsistent with a thermal origin. The third component peaks at $M_{i}\sim-19$ at $\gtrsim 16$ rest-frame days post-FXT, and is compatible with an emerging supernova. We fit the $riz$-band data with a series of power laws and find that the decaying components are in agreement with gamma-ray burst afterglow models, and that the re-brightening may originate from refreshed shocks. By considering EP 240414a in context with all previously reported known-redshift FXT events, we propose that Einstein Probe FXT discoveries may all result from high-redshift gamma-ray bursts, and thus are distinct from the previously discovered lower redshift, lower luminosity population of FXTs.
@@ 1 Introduction
The field of fast X-ray transients (FXTs) is undergoing a revolution since the launch and operation of the Einstein Probe (EP; Yuan et al., 2022) in early 2024. The wide field of view (3600 deg2 instantaneous) allows the sky to be surveyed multiple times in a 24 hr period with the wide-field X-ray Telescope onboard the Einstein Probe (EP-WXT). Its lobster-eye technology typically provides a localisation of 3 arcmin radius in the soft X-ray band of $0.5-4$ keV (e.g., Zhang et al., 2024a, b), with a sensitivity of $1.2\times 10^{-10}$ erg s-1 cm-2 for an exposure of 100 s (Yuan et al., 2022). The survey power is illustrated by the grasp of the EP-WXT (effective area $\times$ field-of-view) being a factor of 10 greater than XMM-Newton, ROSAT and eROSITA (see figure 10 of Yuan et al., 2022). EP’s follow-up X-ray telescope (EP-FXT) can be triggered to follow up and improve the localisation of sources from the EP-WXT to a much smaller radius of about 10 arcsec, while also being significantly more sensitive.
@@
2 Optical counterpart discovery and follow-up
2.1 Discovery of the optical counterpart with LOT
@@
This source was confirmed as the likely counterpart with the 2.56-m Nordic Optical Telescope at 12.29 hrs after the EP trigger (Xu et al., 2024), and a spectrum taken at approximately +15 hr by Jonker et al. (2024b) indicated a featureless continuum and a redshift of $z=0.41$ for the most likely host galaxy SDSS J124601.99-094309.3. We discuss the host galaxy and redshift further in Section 2.3.
2.2 Optical photometric follow-up
@@
2.3 Spectroscopic observation with Gemini and host redshift measurement
With a Gemini-South Director’s Discretionary Time proposal (PI: S. J. Smartt) we were awarded time for longslit spectroscopy of the transient and its host galaxy. However, due to weather and instrument availability constraints we could only observe once the transient had faded from detectability. We centered the slit of the Gemini Multi-Object Spectrograph (GMOS-S) on the host galaxy SDSS J124601.99-094309.3, with a slit angle of 95.2 East of North. This was to capture the position of AT 2024gsa at a separation of $\farcs$9 (see Figure 1) from the host. We obtained $4\times 1800$ s exposures on 17 July 2024 (commencing at 23:17:27 UT/MJD 60508.970) with the R400 grism, the GG455 blocking filter and $\farcs$0 slit width. Dithering was employed in both the spatial ($\farcs$0, for sky subtraction) and wavelength (200 Å, to cover the chip gaps) directions.
@@ 3 Light curve analysis
The optical light curve of AT 2024gsa is shown in Figure 2. The shaded regions highlight distinct phases in the light curve evolution. Phase A represents the rapidly declining early light curve lasting $\sim 1.5$ rest-frame days post-FXT. AT 2024gsa was discovered a few hours following the FXT (see Section 2.1) with a luminosity of $M_{r}\approx-20$ (AB mag).
@@
Phase B (spanning $\sim 1.5-10$ rest-frame days) highlights an abrupt, rapid and achromatic rise in flux by a factor of $\gtrsim 3$, followed by a fast decline. The rapid timescale ($\sim 2$ d rise) and absolute magnitude ($\sim-21$ AB mag) during the peak of phase B bears resemblance to the observed characteristics of luminous fast blue optical transients (LFBOTs; e.g., Perley et al., 2019). However, the red colour of AT 2024gsa during phase B’s peak ($r-z\approx 0.4\pm 0.1$; observer frame) is at odds with LFBOTs. The middle panels in Figure 2 show light curve comparisons between AT 2018cow (Prentice et al., 2018) and AT 2024gsa. Given the redshift of the likely host galaxy for AT 2024gsa ($z=0.4018$; see Section 2.3), we plot the observed $riz$ light curves of AT 2024gsa alongside the $gri$ light curves of AT 2018cow, respectively, for a more meaningful comparison. Although the timescales for the $riz$ ($gri$) light curves for AT 2024gsa (AT 2018cow) are similar, AT 2024gsa is significantly more luminous at peak in the redder $iz$ bands, highlighting the intrinsically red colour in stark contrast to AT 2018cow and LFBOTs in general.
@@
Assuming the phase B peak is thermally driven, the absolute magnitude of $M_{r}\approx-21$ would imply a luminosity of $L\simeq 8\times 10^{43}\,$erg s-1 (with zero bolometric correction). We fit the $riz$-band photometry during phase B with the SuperBol code (Nicholl, 2018). SuperBol converts the extinction-corrected broadband photometry into monochromatic fluxes and performs a blackbody fit (redshift effects are accounted for). We infer a blackbody temperature, $T_{\rm bb}\approx 8000\pm 2000$ K for AT 2024gsa around phase B peak. The large uncertainty on the temperature is due to the fact that the blackbody fit is performed on only three optical bands. The inferred luminosity around phase B peak, integrated across just the $gri$-bands, is $L_{gri}\approx 2.5\times 10^{43}$ erg s-1, whereas the total luminosity integrated over the extrapolated blackbody fit is $L_{\rm bb}\approx 8.5\times 10^{43}$ erg s-1. The fit yields a photospheric radius, $R_{\rm bb}\approx 5\times 10^{15}$ cm. This radius is very large for a compact progenitor origin, thus requiring an extreme ejecta velocity ($v_{\rm ej}\sim 0.7$ c), given the rapid timescale of $\sim 3$ days to (phase B) peak. Although the radius may be plausible for an extended stellar envelope or CSM distribution, as alluded to earlier, the combination of luminous X-ray emission and red optical colours seem to rule out a thermal emission mechanism. This is further illustrated by the requirement of $E_{\rm kin}\gtrsim 4\times 10^{52}$ erg to eject just 0.1 M at this velocity. Thermal emission models for AT 2018cow and other LFBOTs can reproduce the observed light curves as they are extremely hot ($T_{\rm bb}\simeq 30$,000 K) at peak. The red colour of AT 2024gsa makes this explanation untenable for the observed peak at $\sim 3$ days.
@@ 4.1 The nature of EP 240414a/AT 2024gsa
Figure 4: Top panel: Power law $+$ broken power law fits to the full $riz$-band data for AT 2024gsa. Middle panel: Power law $+$ broken power law fits to the $<10$ d $riz$-band data of AT 2024gsa. Bottom panel: Optical-to-X-ray spectrum of the broadband counterpart of EP 240414a, containing data obtained between $4-5$ days post-FXT trigger.
@@
First, we examine the decay rate measured in the first two days. We find that the initial decay rate is consistent across both bands ($\alpha_{1}=-0.31\pm 0.02$ in $r$-band, $\alpha_{1}=-0.33\pm 0.05$ in $i$-band). We find that this observed decay rate is similar to a forward shock in a fast-cooling, optically thin regime ($\nu_{sa}<\nu_{c}<\nu_{obs}<\nu_{m}$): $t^{-0.25}$ and is independent of the circumburst density profile. Here, $\nu_{sa}$, $\nu_{c}$, $\nu_{obs}$ and $\nu_{m}$ are the self-absorption, cooling, observing and characteristic electron break frequencies, respectively.
@@
If we do not include the late time data points, we find that the post-peak decay rates are even steeper: $\alpha_{3}\approx-2.6$, $-2.2\pm 0.4$ and $-2.5\pm 0.1$ at $r$-, $i$- and $z$-band, respectively. These decay rates are consistent with $\alpha_{3}=-p$, where $p$ is the same as that derived from the spectral index. In the GRB scenario this is interpreted as an optically thin, post-jet break decay rate ($\nu_{sa}<\nu_{m}<\nu_{obs}<\nu_{c}$; Sari et al., 1999). A jet break occurs when the jet has decelerated sufficiently such that the beaming cone encompasses the whole shock front, so the observer sees the intrinsic decay rate of the afterglow. We note that the steeper decay rates are consistent with those found by Bright et al. (2024b) with ATCA on timescales of tens of days.
@@
Despite the lack of consistency with canonical GRB afterglow models, the observed variability is not uncommon in optical afterglows of GRBs. Liang et al. (2013) found a number of GRB afterglows that exhibit optical re-brightening on timescales of $\lesssim 1$ day, and a previous Einstein Probe source (EP 240315a) displayed a significant optical and X-ray re-brightening $\approx 3$ days post-burst (Liu et al., 2024c). The most likely explanation is that the optical re-brightening is a consequence of refreshed shocks (Rees & Mészáros, 1998; Sari & Mészáros, 2000) which result from slower shells ejected from the source after the main burst catching up with and re-energising the main afterglow shock at late times, as has been suggested for GRB 030329 (Granot et al., 2003; Moss et al., 2023), GRB 160821B (Lamb et al., 2019), and GRB 170817A (Lamb et al., 2020). We find that the timescale post-FXT and duration of the flare from AT 2024gsa is very similar to that observed by Moss et al. (2023). Other plausible mechanisms include external density fluctuations (e.g., Dai & Lu, 2002), although this is difficult to reconcile with such a sharp flux increase (Nakar & Piran, 2003), or possibly magnetar central engine activity (Gao et al., 2015). Unfortunately, we do not have comprehensive spectral and temporal coverage to rule out these other scenarios with confidence.
@@ 4.2 EP 240414a in context with other FXTs
EP has released GCNs for 27 FXT events, with eight optical counterparts discovered, and redshifts measured for just five of these. Generally, they appear to be significantly more distant than the redshifts inferred for the XMM-Newton (Alp & Larsson, 2020) and Chandra (Quirola-Vásquez et al., 2022, 2023) FXTs. To determine whether EP is detecting higher redshift FXT analogues to the XMM-Newton and Chandra samples, or whether it is instead uncovering a distinct population of luminous, high-redshift FXTs, we compare the various samples of FXT events that have reported redshift values.
@@
To this set, we add XT 030206, XT 070618, XT 100424, XT 110621, XT 151128, XT 151219 and XT 161028 extracted from the XMM-Newton catalogue presented by Alp & Larsson (2020) (hereafter referred to as the A&L2020 sample). From archival Chandra observations spanning $2000-2014$, we include XRT 041230 and XRT 080819 (Quirola-Vásquez et al., 2022, hereafter the QV+2022 sample), and from Chandra observations spanning $2014-2022$, we include XRT 150322 (Xue et al., 2019b; Quirola-Vásquez et al., 2023), XRT 161125 (Quirola-Vásquez et al., 2023), XRT 170901 (Lin et al., 2019; Quirola-Vásquez et al., 2023), XRT 191223 (Quirola-Vásquez et al., 2023) and XRT 210423 (Lin et al., 2021; Quirola-Vásquez et al., 2023), hereafter the QV+2023 sample.
@@
Given the EP-WXT observing strategy of $t_{\rm exp}=900-1200$ s (Yuan et al., 2022), a source with $L_{\rm iso}=10^{45}$ erg s-1 would only be detectable at a redshift, $z\lesssim 0.1$. Hence, the X-ray luminosity parameter space probed by EP is distinct from that which most previously known FXTs occupy. EP is designed such that it has a higher luminosity threshold than XMM-Newton and Chandra, making it insensitive to the parameter space occupied by the historical FXT samples (see Figure 5 for a visualisation of the sensitivity limits of EP, XMM-Newton and Chandra). Conversely, XMM-Newton and Chandra have not detected FXTs that match those uncovered by EP; this implies that EP-detected events are intrinsically (volumetrically) rare events that have been routinely missed by the narrow fields of view of these observatories. However, the large field of view of EP has ensured their detection. Thus, it appears that EP is primarily sensitive to the rare, luminous events that are broadly consistent with cosmological GRB events, whereas the deep, narrower field-of-view observatories are more suited to uncovering intrinsically fainter, but volumetrically more common, FXTs.
The one exception to this broad picture is XRT 161125, which has $L_{\rm iso}=2\times 10^{47}$ erg s-1, and a redshift, $z=0.35^{+0.05}_{-0.15}$ (Quirola-Vásquez et al., 2023). This redshift is based on an unconfirmed association with the marginally resolved $r=24.2$ mag host galaxy in the DECam Legacy Surveys (Dey et al., 2019). Thus, the association is far from secure and it is plausible that this is merely a chance coincidence (Quirola-Vásquez et al. 2023 estimate a chance association probability of $<0.095$), with XRT 161125 being at a much larger distance, making it similar in physical nature to the high-redshift EP transient sample. In Figure 5, we illustrate where in the luminosity/redshift parameter space XRT 161125 would lie for a range of redshifts ($z\geq 0.2$). We find that the isotropic luminosity inferred for XRT 161125 closely resembles that of the four EP FXTs at $z\gtrsim 1.5$ (EP 240315a, EP 240801a, EP 240804a and EP 240806a).
Alternatively, if the host galaxy association for XRT 161125 is correct, then we note that XRT 161125 possesses a (projected) offset of $\approx 13.2$ kpc (Quirola-Vásquez et al., 2023). This value is approximately half the separation observed here for EP 240414a/AT 2024gsa (27 kpc; see Section 2.3).
This comparison presented here, between the XMM-Newton, Chandra and EP FXT populations, tentatively suggests that the known FXTs sit in two distinct populations, namely: (i) a high-redshift set of X-ray emitters that are likely all (or mostly) the prompt X-ray emission of GRBs, and (ii) the low redshift, $L_{\rm X}\simeq{\rm a\ few}\times 10^{42}-10^{46}$ erg s-1 population discovered by Alp & Larsson (2020) and Quirola-Vásquez et al. (2022, 2023).

### v2

# Identification of the optical counterpart of the fast X-ray transient EP240414a

@@ Abstract
Fast X-ray transients (FXTs) are extragalactic bursts of X-rays first identified in archival X-ray data, and now routinely discovered by the Einstein Probe in real time, which is continuously surveying the night sky in the soft ($0.5-4$ keV) X-ray regime. In this Letter, we report the discovery of the second optical counterpart (AT 2024gsa) to an FXT (EP240414a). EP240414a is located at a projected radial separation of 27 kpc from its likely host galaxy at $z=0.4018\pm 0.0010$. The optical light curve of AT 2024gsa displays three distinct components. The initial decay from our first observation is followed by a re-brightening episode, displaying a rapid rise in luminosity to an absolute magnitude of $M_{r}\sim-21$ after two rest-frame days. While the early optical luminosity and decline rate is similar to luminous fast blue optical transients, the colour temperature of AT 2024gsa is distinctly red and we show that the peak flux is inconsistent with a thermal origin. The third component peaks at $M_{i}\sim-19$ at $\gtrsim 16$ rest-frame days post-FXT, and is compatible with an emerging supernova. We fit the $riz$-band data with a series of power laws and find that the decaying components are in agreement with gamma-ray burst afterglow models, and that the re-brightening may originate from refreshed shocks. By considering EP240414a in context with all previously reported known-redshift FXT events, we propose that Einstein Probe FXT discoveries may predominantly result from (high-redshift) gamma-ray bursts, and thus appear to be distinct from the previously discovered lower redshift, lower luminosity population of FXTs.
@@ 1 Introduction
Fast X-ray transients (FXTs) are extragalactic, short duration bursts of X-rays lasting a few seconds to thousands of seconds, typically not associated with a higher (i.e., gamma-ray) or lower (i.e., optical, radio) energy counterpart. The field of FXTs is undergoing a revolution since the launch and operation of the Einstein Probe (EP; Yuan et al., 2022) in early 2024. The wide field of view (3600 deg2 instantaneous) allows the sky to be surveyed multiple times in a 24 hr period with the wide-field X-ray Telescope onboard the Einstein Probe (EP-WXT). Its lobster-eye technology typically provides a localisation of 3 arcmin radius in the soft X-ray band of $0.5-4$ keV (e.g., Zhang et al., 2024a, b), with a sensitivity of $1.2\times 10^{-10}$ erg s-1 cm-2 for an exposure of 100 s (Yuan et al., 2022). The survey power is illustrated by the grasp of the EP-WXT (effective area $\times$ field-of-view) being a factor of 10 greater than XMM-Newton, ROSAT and eROSITA (see figure 10 of Yuan et al., 2022). EP’s follow-up X-ray telescope (EP-FXT) can be triggered to follow up and improve the localisation of sources from the EP-WXT to a much smaller radius of about 10 arcsec, while also being significantly more sensitive.
@@
2 Discovery of the optical counterpart with LOT
@@
This source was confirmed as the likely counterpart with the 2.56-m Nordic Optical Telescope at 12.29 hrs after the EP trigger (Xu et al., 2024), and a spectrum taken at approximately +15 hrs by Jonker et al. (2024b) indicated a featureless continuum and a redshift of $z=0.41$ for the most likely host galaxy SDSS J124601.99-094309.3.
Following discovery, we launched a multi-wavelength observing campaign to sample the evolution of the optical counterpart; for details of our observing campaign, see Appendix A. Our optical light curve is presented in Figure 3 and the summary of photometric observations is provided in Table 1. We also present the Swift/XRT data and use it to constrain the optical to X-ray spectral energy distribution.
@@
3 Spectroscopic observation with Gemini and host redshift measurement
With a Gemini-South Director’s Discretionary Time proposal (PI: S. J. Smartt) we were awarded time for longslit spectroscopy of the transient and its host galaxy. However, due to weather and instrument availability constraints we could only observe once the transient had faded from detectability. We centred the slit of the Gemini Multi-Object Spectrograph (GMOS-S) on the host galaxy SDSS J124601.99-094309.3, with a slit angle of 95.2 East of North. This was to capture the position of AT 2024gsa at a separation of $\farcs$9 (see Figure 1) from the host. We obtained $4\times 1800$ s exposures on 17 July 2024 (commencing at 23:17:27 UT/MJD 60508.970) with the R400 grism, the GG455 blocking filter and $\farcs$0 slit width. Dithering was employed in both the spatial ($\farcs$0, for sky subtraction) and wavelength (200 Å, to cover the chip gaps) directions.
@@ 4 Light curve analysis
The optical light curve of AT 2024gsa is shown in Figure 3. The shaded regions highlight distinct phases in the light curve evolution. Phase A represents the rapidly declining early light curve lasting $\sim 1.5$ rest-frame days post-FXT. AT 2024gsa was discovered a few hours following the FXT (see Section 2) with a luminosity of $M_{r}\approx-20$ (AB mag).
@@
Phase B (spanning $\sim 1.5-10$ rest-frame days) highlights an abrupt, rapid and achromatic rise in flux by a factor of $\gtrsim 3$, followed by a fast decline. The rapid timescale ($\sim 2$ d rise) and absolute magnitude ($\sim-21$ AB mag) during the peak of phase B bears resemblance to the observed characteristics of luminous fast blue optical transients (LFBOTs; e.g., Perley et al., 2019). However, the red colour of AT 2024gsa during phase B’s peak ($r-z\approx 0.4\pm 0.1$; observer frame) is at odds with LFBOTs. The middle panels in Figure 3 show light curve comparisons between AT 2018cow (Prentice et al., 2018) and AT 2024gsa. Given the redshift of the likely host galaxy for AT 2024gsa ($z=0.4018$; see Section 3), we plot the observed $riz$ light curves of AT 2024gsa alongside the $gri$ light curves of AT 2018cow, respectively, for a more meaningful comparison. Although the timescales for the $riz$ ($gri$) light curves for AT 2024gsa (AT 2018cow) are similar, AT 2024gsa is significantly more luminous at peak in the redder $iz$ bands, highlighting the intrinsically red colour in stark contrast to AT 2018cow and LFBOTs in general.
@@
Assuming the phase B peak is thermally driven, the absolute magnitude of $M_{r}\approx-21$ would imply a luminosity of $L\simeq 8\times 10^{43}\,$erg s-1 (with zero bolometric correction). We fit the $riz$-band photometry during phase B with the SuperBol code (Nicholl, 2018). SuperBol converts the extinction-corrected broadband photometry into monochromatic fluxes and performs a blackbody fit (redshift effects are accounted for). We infer a blackbody temperature, $T_{\rm bb}\approx 8000\pm 2000$ K for AT 2024gsa around phase B peak. The large uncertainty on the temperature is due to the fact that the blackbody fit is performed on only three optical bands. The inferred luminosity around phase B peak, integrated across just the $gri$-bands, is $L_{gri}\approx 2.5\times 10^{43}$ erg s-1, whereas the total luminosity integrated over the extrapolated blackbody fit is $L_{\rm bb}\approx 8.5\times 10^{43}$ erg s-1. The fit yields a photospheric radius, $R_{\rm bb}\approx 5\times 10^{15}$ cm. This radius is very large for a compact progenitor origin, thus requiring an extreme ejecta velocity ($v_{\rm ej}\sim 0.7$ c), given the rapid timescale of $\sim 3$ days to (phase B) peak. Although the radius may be plausible for an extended stellar envelope or CSM distribution, as alluded to earlier, the combination of luminous X-ray emission and red optical colours seem to rule out a thermal emission mechanism. This is further illustrated by the requirement of $E_{\rm kin}\gtrsim 4\times 10^{52}$ erg to eject just 0.1 M at this velocity. Thermal emission models for AT 2018cow and other LFBOTs can reproduce the observed light curves as they are extremely hot ($T_{\rm bb}\simeq 30$,000 K) at peak. The red colour of AT 2024gsa makes this explanation untenable for the observed peak at $\sim 3$ days. We present a color comparison for AT 2024gsa with AT 2018cow, GRB 071010A and SN 1998bw in Appendix B and Figure 6. This illustrates the red color of AT2024gsa, and while the early color evolution may be similar to SN 1998bw, the very high luminosity requires a large photospheric radius and thus an extreme ejecta velocity, as described above.
@@ 5.1 The nature of EP240414a/AT 2024gsa
Figure 4: Top panel: Power law $+$ broken power law fits to the full $riz$-band data for AT 2024gsa. Temporal indices during the initial decay around phase A are $\alpha_{1}=-0.31\pm 0.02$, $-0.33\pm 0.05$ for $r$ and $i$-band respectively. The indices for the post-peak decline are $\alpha_{3}=-1.4\pm 0.4$, $-1.04\pm 0.07$ and $-0.93\pm 0.09$ for $r$, $i$ and $z$-band respectively.
Middle panel: Power law $+$ broken power law fits to the $<10$ d $riz$-band data of AT 2024gsa. The indices for the post-peak decline are steeper after excluding the phase C component, with $\alpha_{3}=-2.6\pm 0.4$, $-2.2\pm 0.4$ and $-2.5\pm 0.1$ for $r$, $i$ and $z$-band, respectively. Bottom panel: Optical-to-X-ray spectrum of the broadband counterpart of EP240414a, containing data obtained between $4-5$ days post-FXT trigger.
@@
First, we examine the decay rate measured in the first two days. We find that the initial decay rate is consistent across both bands ($\alpha_{1}=-0.31\pm 0.02$ in $r$-band, $\alpha_{1}=-0.33\pm 0.05$ in $i$-band). We find that this observed decay rate is similar to a forward shock in a fast-cooling, optically thin regime ($\nu_{sa}<\nu_{c}<\nu_{obs}<\nu_{m}$): $t^{-0.25}$ and is independent of the circum-burst density profile. Here, $\nu_{sa}$, $\nu_{c}$, $\nu_{obs}$ and $\nu_{m}$ are the self-absorption, cooling, observing and characteristic electron break frequencies, respectively.
@@
If we do not include the late time data points, we find that the post-peak decay rates are even steeper: $\alpha_{3}\approx-2.6\pm 0.4$, $-2.2\pm 0.4$ and $-2.5\pm 0.1$ at $r$-, $i$- and $z$-band, respectively. These decay rates are consistent with $\alpha_{3}=-p$, where $p$ is the same as that derived from the spectral index. In the GRB scenario this is interpreted as an optically thin, post-jet break decay rate ($\nu_{sa}<\nu_{m}<\nu_{obs}<\nu_{c}$; Sari et al., 1999). A jet break occurs when the jet has decelerated sufficiently such that the beaming cone encompasses the whole shock front, so the observer sees the intrinsic decay rate of the afterglow. We note that the steeper decay rates are consistent with those found by Bright et al. (2024b) with ATCA on timescales of tens of days.
@@
Despite the lack of consistency with canonical GRB afterglow models, the observed variability is not uncommon in optical afterglows of GRBs. Liang et al. (2013) found a number of GRB afterglows that exhibit optical re-brightening on timescales of $\lesssim 1$ day, and a previous Einstein Probe source (EP240315a) displayed a significant optical and X-ray re-brightening $\approx 3$ days post-burst (Liu et al., 2024c). The most likely explanation is that the optical re-brightening is a consequence of refreshed shocks (Rees & Mészáros, 1998; Sari & Mészáros, 2000) which result from slower shells ejected from the source after the main burst catching up with and re-energising the main afterglow shock at late times, as has been suggested for GRB 030329 (Granot et al., 2003; Moss et al., 2023), GRB 160821B (Lamb et al., 2019), and GRB 170817A (Lamb et al., 2020). The dynamical time scale in the ejecta is comparable to the time of observation within a decelerating blastwave picture (Moss et al., 2023). We note that the rebrightening episode during phase B starts at $t\sim 1.5$ rest-frame days with a rise time of $\Delta t\sim 1.5$ rest-frame days, thus satisfying $\Delta t/t\sim 1$. We find that the timescale post-FXT and duration of the flare from AT 2024gsa is very similar to that observed by Moss et al. (2023). Other plausible mechanisms include external density fluctuations (e.g., Dai & Lu, 2002), although this is difficult to reconcile with such a sharp flux increase (Nakar & Piran, 2003), or possibly magnetar central engine activity (Gao et al., 2015). Unfortunately, we do not have comprehensive spectral and temporal coverage to rule out these other scenarios with confidence.
@@ 5.2 EP240414a in context with other FXTs
EP has released GCNs for 27 FXT events, with eight optical counterparts discovered, and redshifts measured for just five of these. [^2] As of 14 August 2024. Generally, they appear to be significantly more distant than the redshifts inferred for the XMM-Newton (Alp & Larsson, 2020) and Chandra (Quirola-Vásquez et al., 2022, 2023) FXTs. Here we qualitatively compare the properties of the various samples of FXT objects, to explore their similarities and differences. Despite the properties and discovery methods of the Chandra and XMM-Newton FXT samples being distinct from those uncovered by EP, a comparison of their properties can still provide useful insights, as presented below.
@@
To this set, we add XT 030206, XT 070618, XT 100424, XT 110621, XT 151128, XT 151219 and XT 161028 extracted from the XMM-Newton catalogue presented by Alp & Larsson (2020) (hereafter the A&L2020 sample). From archival Chandra observations spanning $2000-2014$, we include XRT 041230 and XRT 080819 (Quirola-Vásquez et al., 2022, hereafter the QV+2022 sample), and from Chandra observations spanning $2014-2022$, we include XRT 150322 (Xue et al., 2019b; Quirola-Vásquez et al., 2023), XRT 161125 (Quirola-Vásquez et al., 2023), XRT 170901 (Lin et al., 2019; Quirola-Vásquez et al., 2023), XRT 191223 (Quirola-Vásquez et al., 2023) and XRT 210423 (Lin et al., 2021; Quirola-Vásquez et al., 2023), hereafter the QV+2023 sample.
@@
Given the EP-WXT observing strategy of $t_{\rm exp}=900-1200$ s (Yuan et al., 2022), a source with $L_{\rm iso}=10^{45}$ erg s-1 would only be detectable at a redshift, $z\lesssim 0.1$. Hence, the X-ray luminosity parameter space probed by EP is distinct from that which most previously known FXTs occupy. EP is designed such that it has a higher luminosity threshold than XMM-Newton and Chandra, making it insensitive to the parameter space occupied by the historical FXT samples (see Figure 5 for a visualisation of the sensitivity limits of EP, XMM-Newton and Chandra). Conversely, XMM-Newton and Chandra have not detected FXTs that match those uncovered by EP; this is indicative of EP-detected events being intrinsically (volumetrically) rare events that have been routinely missed by the narrow fields of view of these observatories. However, the large field of view of EP has ensured their detection. Thus, it appears that EP is primarily sensitive to the rare, luminous events that are broadly consistent with cosmological GRB events (out to high-$z$), whereas the deep, narrower field-of-view observatories are more suited to uncovering intrinsically fainter (but volumetrically more common) FXTs.
The one exception to this broad picture is XRT 161125, which has $L_{\rm iso}=2\times 10^{47}$ erg s-1, and a redshift, $z=0.35^{+0.05}_{-0.15}$ (Quirola-Vásquez et al., 2023). This redshift is based on an unconfirmed association with the marginally resolved $r=24.2$ mag host galaxy in the DECam Legacy Surveys (Dey et al., 2019). Thus, the association is far from secure and it is plausible that this is merely a chance coincidence (Quirola-Vásquez et al. 2023 estimate a chance association probability of $<0.095$), with XRT 161125 being at a much larger distance, making it similar in physical nature to the high-redshift EP transient sample. In Figure 5, we illustrate where in the luminosity/redshift parameter space XRT 161125 would lie for a range of redshifts ($z\geq 0.2$). We find that the isotropic peak luminosity inferred for XRT 161125 closely resembles that of the four EP FXTs at $z\gtrsim 1.5$ (EP240315a, EP240801a, EP240804a and EP240806a).
Alternatively, if the host galaxy association for XRT 161125 is correct, then we note that XRT 161125 possesses a (projected) offset of $\approx 13.2$ kpc (Quirola-Vásquez et al., 2023). This value is approximately half the separation observed here for EP240414a/AT 2024gsa (27 kpc; see Section 3).